OpenStudy (anonymous):

A computer repair company charges a fee and an hourly charge. After two hours the repair bill is $110, and after three hours it is $150. How much would it cost for 1.5 hours of work?

OpenStudy (anonymous):

h=hourly rate 2h=110 3h=150 3h-2h=150-110 therefore, the hourly rate is 40 and hour. for two hours they charged 110, so 40*2=80. 110-80 is 30. therefore, the beginning fee they charge is 30. It is 30 to begin with then 40 for each hour of work.

OpenStudy (anonymous):

Ok so for the fee and the cost of 1.5 hours would be 30+40+20? $90?